# Prompts to Summaries: Zero-Shot Language-Guided Video Summarization

### v1

@@ 4.1.4 Results
Supervised methods
re-SEQ2SEQ [21] - 63.9
CLIP-It [20] 54.2 66.3
MAVS  [28] - 67.5
iPTNet [19] 54.5 63.4
A2Summ [18] 55.0 63.4
PGL-SUM [29] 57.1 -
Unsupervised methods
AC-SUM-GAN [17] 50.8 60.6
RS-SUM [16] 52.0 61.4
SegSum [15] 54.07 62.02
Zero-shot methods
Random 44.89 56.43
Ours 56.73 62.21
Performance comparison. Although our approach performs competitively with supervised methods, our primary focus is to compare the performance of our model with previous unsupervised methods. Notably, our approach is also data-free and support user guidelines via text.
The comparative analysis is presented in Table 2. The results demonstrate that our unsupervised zero-shot model surpasses all competing unsupervised (data-hungry) methods. Specifically, our approach achieves a 2.7% improvement in the sumMe dataset and a 0.2% improvement in the TVSum dataset compared to the state-of-the-art method [15], which incorporates the temporal video segmentation and summarization process from KTS [30] and the CA-SUM architecture [16]. Also we improve by 4.7% in the sumMe dataset and 0.81% in the TVSum dataset upon [16] which leverages a powerful self-supervised learning method recently employed for training language models. All that and our method requires no training data at all.
Overall, this suggests that VideoLMs and LLMs can be utilized and repurposed for the video summarization task without the need for any previous knowledge about user annotations or data distribution .
(a) Visualization of the video summary for the video titled "Bike Polo" utilizing our model from SumMe dataset
(b) Visualization of the video summary for the video titled "4wU_LUjG5Ic" utilizing our model from TVSum dataset
Figure 5: Visualization of the selections made by the our method, user summaries, and KTS segments assigned to each segment in videos from the SumMe and TVSum datasets. The blue bars represent human frame selections for the video, with each row indicating a human summary. The yellow bars depicts the video fragments that will be art of the final summary , marked by black lines. The green bar illustrates our model’s frame selection for the summary, while the red bar represents frames that are not selected. Additionally, video frames from each selected shot in our summary are included for better visualization.
4.1.5 Qualitative example
First we introduce the Kernel Temporal Segmentation (KTS) [30], a methods used to divide the video into coherent temporal segments based on changes in content over time, it is often used as labels for training to other video summarization methods. To visually showcase the temporal selection process of our model for video summarization, we present the selections made by our model (in red), user summaries (blue), and the ground truth KTS segments (yellow) across videos from the SumMe and TVSum datasets, as shown in Figure 5.
In Figure 5(a), which illustrates the “Bike Polo” video from the SumMe dataset, we observe that user summaries do not align well with the predefined KTS segments. This misalignment makes it challenging to identify the most optimal segments for constructing an accurate summary. However, our model successfully selects KTS segments that are most aligned with user preferences—specifically, segments 14 and 15.
In contrast, in video “4wU_LUjG5Ic” from the TVSum dataset, shown in Figure 5(b), the user summaries match both the KTS segmentation and ours. Nevertheless, we notice that the final (single) segment selected by our model—segment 26—is not the most aligned with user choices. Segments like 1 or 3 may have been more representative. This discrepancy is likely due to constraints imposed by the final summary length or limitations in the generated textual description, which may have failed to sufficiently highlight the importance of these segments, leading to lower relevance scores from the LLM.
4.2 Query-Focused video summarization
4.2.1 Datasets
For evaluating the performance of our Query-focused video summarization approach, we utilize QFVS dataset  [27], which consists of 4 long videos based on existing UT Egocentric (UTE) dataset  [31], each video is 3-5 hours long, paired with 45-46 user-queries and user-annotated summaries, containing a diverse set of events. Each video was annotated by 3 users, the video was segmented into 5-second non-overlapping shots, each with captions that are present in it like SKY, STREET, ROAD and CAR. The 3 users captions where aggregated, per video-query, into one caption-dense annotations, and then used to obtain video-query summaries using greedy algorithm. Those aggregated summaries, Oracle summaries, achieve better agreements between the user than the inter-user consensus.
4.2.2 Evaluation protocol
Method name Property Doesn’t require training data Vid 1 Vid 2 Vid 3 Vid 4 Avg
SeqDPP [24] Supervised 36.59 43.67 25.26 18.15 30.92
SH-DPP [23] Supervised 35.67 42.74 36.51 18.62 33.38
QFVS [27] Supervised 48.68 41.66 56.47 29.96 44.19
CLIP-It : ResNet [20] Supervised 55.19 51.03 64.26 39.47 52.49
CLIP-It : CLIP-Image [20] Supervised 57.13 53.60 66.08 41.41 54.44
Ours : MinMax norm Zero-Shot 51.47 50.86 61.29 37.51 50.28
Ours : Exp norm Zero-Shot 48.66 50.00 55.26 42.63 49.14
Ours : (MinMax + Exp) norm Zero-Shot 53.57 55.66 63.25 41.20 53.42
4.2.3 Quantitative results
In Table 3, we compare the F1 scores of our method against five supervised baselines: Seq-DPP [24], SH-DPP [23], QFVS [27], CLIP-It: Resnset [20], and CLIP-It: Clip Image [20]. Following the evaluation protocol of [27], which conducts four rounds of experiments by leaving out one video for testing and one for validation while using the remaining two for training, we evaluate our model—despite it being training-free—on four separate splits, each containing a single test video. We then average the F1 scores across these splits to report the final performance.
As shown in Table 3, our training-free, zero-shot approach consistently outperforms the four classical supervised baselines on the QFVS dataset. The only method outperforming us slightly is the data hungry CLIP-It [20] which requires and per frame exhaustive annotation and is highly sensitive to the image backbone it fine-tunes: when CLIP-It relies on ResNet-50 [42], our method leads by +1 pp F1—despite using no task-specific data. For a like-for-like comparison we also adopt the stronger CLIP embeddings [36]; in this setting our zero-shot model that reached an F1 of 53.42 %, virtually matching the 54.44 % achieved by the fully supervised CLIP-It. A per-video analysis further underlines our robustness: we surpass CLIP-It by +2 pp on Video 2 and equal its performance on the notoriously challenging Video 4. Altogether, these results demonstrate that our method delivers state-of-the-art accuracy without any fine-tuning, hyper-parameter search, or prior exposure to the dataset, making it both effective and readily deployable.
Further analysis. Striving for optimal performance in our model, we evaluated the effects of different normalization techniques : Min-Max normalization, Exponential normalization, and a combined approach. The combined method was introduced based on its strong performance on the SumMe and TVSum datasets, aiming to benefit from both scale consistency and enhanced contrast. As predicted, the combination of both normalization methods yielded the best results across all 4 videos, confirming its ability to balance sensitivity and robustness in representing importance scores. This comparison highlights the combined method as the most effective approach for video summarization in this context, as shown in Table 3. Notably, all of our variants (normalization techniques) consistently achieved better results than all of our competing methods except of Clip-it.
Figure 6: Result of our method on the QFVS dataset. The middle columns framed by black outline, showcase frames sampled from video 3, 3 hours in duration, from QFVS dataset. Summary 1 shows sampled frames for our summary, given the query "Focus on scenes containing Chair and/or Tree". Summary 2 shows sampled frames for our summary, given the query "Focus on scenes containing Food and/or Hands". Green boxes highlight the entities specified in the respective queries.
4.2.4 Qualitative example
To demonstrate our model’s ability to guide the summarization process using different user queries expressed in natural language, we present (qualitative results in) Figure 6. Specifically, as an example, we apply our model to an input video (Video 3) from the QFVS dataset—a 3-hour video recorded using a head-mounted camera. The video captures a wide range of daily activities, including driving to various locations, shopping at a supermarket, and preparing and eating food. Summary 1 and Summary 2 show sampled frames from the summaries generated by our model in response to two different queries: (i) ’Focus on scenes containing Chair and / or Tree’ (ii) ’Focus on scenes containing Food and / or Hands’
These results highlight the model’s ability to assign high importance scores to segments, scenes, or events that align with user-defined preferences. Notably, although the input video remains the same, the generated summaries differ significantly based on the provided query, demonstrating the model’s adaptability to user intent, e.g., for the query “Focus on scenes containing Chair and/or Tree”, as illustrated by sample images from the generated summary, our model successfully selects relevant shots—containing chairs and/or trees (highlighted with green boxes)—that fulfill the intent of the query.
4.3 Fine-grained video summarization
Although QFVS is a useful text-guided benchmark, its queries are mostly are limited to short, literal keywords (e.g. car, street, pets) and therefore do not probe temporal reasoning or long-tail semantics. To evaluate fine-grained, intent-aware summarization, we introduce a new dataset, VidSum-Reason.
4.3.1 VidSum-Reason
Our new benchmark for fine-grained, reasoning-driven video summarization from natural language Queries.
VidSum-Reason consists of 9 videos ranging from 2–6 minutes in duration, covering a diverse range of content including sports, DIY projects, movie trailers, award-winning CGI videos and selected videos from Youtube-8M dataset [43]. Each video is paired with one to three queries, yielding 20 video–query pairs in total. Unlike QFVS—where queries which merely refer to obvious visual entities—VidSum-Reason poses richer, more nuanced prompts that test intent interpretation, temporal reasoning, and the use of external knowledge. VidSum-Reason queries not only prioritize certain content but also serve as filters to exclude scenes, e.g., filtering violence or Excluding PG-13+ material. We group the queries into four classes:
  1. (i)
    Standard: Queries that describe a specific object or action observable in the video, similar to QFVS but potentially multi-word.
  2. (ii)
    Standard with Special Attribute: Queries that reference visual content along with a descriptive attribute, requiring models to focus on both presence and characteristics.
  3. (iii)
    Reasoning: Queries that require some level of temporal or contextual reasoning beyond simple object recognition.
  4. (iv)
    Reasoning with General Knowledge: Queries that rely on external knowledge or a broader understanding of context to identify relevant segments.
This design provides a challenging testbed for models that aim to summarize videos in a genuinely query-focused, knowledge-aware manner.
Figure 7: Comparison of F1-scores between our approach and the random baseline across the four query classes in the  VidSum-Reason dataset. Our model consistently outperforms the random baseline across all classes. The average F1-score is 43.4 for our model and 34.56 for the random baseline, demonstrating the effectiveness of query-guided summarization.
To generate ground truth annotations for each video query pair, the video was segmented into non-overlapping 2-second segments, following the same approach used in the TVSum dataset [26]. Each segment was rated on a scale from 1 (not relevant) to 5 (very relevant) based on how well it matched the content of a specific user query. If a segment was found not relevant to the query, then, it was evaluated based on its overall importance to the main theme, message, or core event of the video, using the same 1–5 scale. This dual scoring approach ensures that both query-relevant and contextually important segments are captured. The resulting ground truth provides a robust, high-quality annotation: it enables evaluation of how well a method can identify content directly related to user intent (query-focused summarization), while also preserving the ability to generate generic video summaries that reflect the video’s core storyline. Annotations were initially created by a single user (author) and then reviewed by two additional individuals to ensure consistency, accuracy, and overall quality.
4.3.2 Evaluation protocol
We utilize the same method commonly employed for evaluating generic video summaries — the keyshot evaluation method — originally introduced by [6] and widely adopted in state-of-the-art video summarization research. This metric evaluates the similarity between the generated summary (A), constrained by a duration budget, and the user-provided summary (B). It computes the temporal overlap between the two using precision (P) and recall (R), as defined in Equations 9 and 10.
To generate a summary from frame-level importance scores, we uniformly divide each video into non-overlapping fragments, where each fragment spans $d\%$ of the video’s total duration. For each fragment, we compute a fragment-level score by averaging the importance scores of all frames within that fragment. To select the final summary, we apply the Knapsack algorithm, subject to a summary budget of $M\%$ of the total duration of the video, as in TVSum and SumMe [6, 26]. In our experiments, we set the fragment size to 3% and the summary budget to 36% (12 fragments). Additional design details are provided in the section C of the Appendix.
4.3.3 Quantitative results
To assess the effectiveness of our fine-grained video summarization model, we conduct a quantitative comparison against a random baseline. As illustrated in Figure 7, we report the total F1 score (obtained by averaging over the splits accuracy) and the F1 scores across four different query categories defined in our  VidSum-Reason dataset: Standard, Standard with Special Attribute, Reasoning, and Reasoning with General Knowledge. Notably, the performance gap is prominent in almost every query type, such as with Reasoning with General Knowledge, where understanding of temporal dependencies and commonsense knowledge is required. Unfortunately, the accuracy of "Stranded with special attributive" query class wasn’t as high as the rest, we suspect, this due to the difficulty level of the query, such as ’focus on bright colors’ which might be hard to interpret.
Figure 8: Visualization of query-guided video summarization. This figure compares the ground-truth and predicted summaries for the query: “Highlight crossover vehicles in the video.” The top plot shows the ground-truth importance scores (in blue) with the user summary overlaid in red boxes, while the bottom plot shows the same scores with the predicted summary in green boxes. Sample video frames are provided for visual reference. Frames common to both summaries that align with the query are highlighted in green; those that represent key events of the video are highlighted in black. This illustrates the model’s ability to capture semantically relevant content aligned with both the query and the video’s main theme.
The random baseline was computed by averaging the F1-scores of 100 randomized tests per query category [6], ensuring a fair and statistically stable comparison. On average, our model achieves a total F1-score of 43.4, while the random baseline yields 34.56, demonstrating the advantage of leveraging text queries for fine-grained summarization. These results validate our model’s ability to adapt to diverse query types and produce contextually relevant summaries that align with user intent.
4.3.4 Qualtitave example
4.4 Ablation
To assess the contribution of each component in our method, we conducted a series of ablation experiments. Specifically, we evaluate the impact of (i) image encoder for calculating frame weights within a scene (Section 3.5.2), (ii) global vs. local visual inputs for description generation (Section 3.2), and (iii) hyperparameter analysis with different model variations.
Image encoder. we investigated the effect of the image encoder on the summarization quality by evaluating DINO model [35] and comparing it to CLIP while utilizing different values of frame consistency weight $\sigma$ and segment duration $W$. We measured their performance on the TVSum and SumMe datasets.
As illustrated in Figures 10 and 10, while some DINO variations achieved competitive results, specifically on TVSum dataset achieving a high 61.5% f1-score, they consistently underperformed compared to CLIP. The CLIP-image encoder produced the highest F1-scores across both datasets, demonstrating stronger semantic representation and robustness in detecting meaningful video segments.
Global vs. Local Visual Inputs for Description Generation. To understand the impact of visual context on the quality and informativeness of scene-level descriptions, we conducted experiments comparing two approaches: global (masking a scene from a video) input versus local-only (single scene) input. The idea is to compare whether it’s easier to assess a scene’s importance by contrasting it with the entire video, or by comparing the full video without that scene to the original full video; i.e., the gain by adding it vs the loss of removing it.
Figure 9: Ablating different Image encoders and Visual context input for description generation. F1 scores results on the test set of SumMe dataset employing MinMax normalization across different segment durations and $\sigma$ values ($\sigma\in[0.0,1.0]$).
Figure 10: Ablating different Image encoders and Visual context input for description generation. F1 scores results on the test set of TVSum dataset employing Exponential (Exp) normalization across different segment durations and $\sigma$ values ($\sigma\in[0.0,1.0]$).
In the global approach, we masked a specific scene by replacing its frames with black frames that includes the text "MASKED SCENE", while keeping the rest of the video intact. This masked video was then passed to the VideoLM with a modified VideoLM’s prompt template that asked the model, not only describe the video in details as demonstrated in 3.2, but specifically to describe if there is a masked section in the video or scene. Similar to VideoLM’s prompt template, the LLMs prompt template was also modified enforcing the LLM to assign an importance score based on the missing information from the global masked textual description on a certain scene. This allowed the model to reason about the scene’s absence in the context of the entire video, implicitly assessing its importance based on the degradation of the global narrative. In contrast, the local approach involves using only the frames of the scene in question as input to generate a description. This method relies purely on the scene’s standalone visual content without incorporating broader temporal context. Contrary to our expectations, the local-only input consistently outperformed the global masked method across both the TVSum and SumMe datasets, as shown in Figure 10 and Figure 10. Interestingly, in SumMe, while the global approach (where the scene in question is masked) achieved competitive results-54.76% f1 score, this can be attributed to the nature of certain SumMe videos, which include more abstract scenes (e.g., burning matchsticks video). These scenes are often difficult to assess using local information alone. Nevertheless, the use of local descriptions proved to be more accurate, focused, and relevant to the content of the scene. In contrast, the global approach occasionally introduced missing information, misleading content, or overgeneralized descriptions due to the absence of visual evidence from the masked segment, leading to noisier importance estimates. These findings suggest that, in the context of description-driven video summarization, local visual input provides a stronger grounding for generating semantically rich and precise descriptions, particularly when the objective is to evaluate the informativeness of individual scenes in isolation.
Normalization function. To further validate our normalization function choices, we conducted additional experiments using different normalization techniques with varying values of frame consistency weights ($\sigma\in[0.0,1.0]$) and segment durations ($W\in\{1,3\}$ seconds). Specifically, we evaluated our three candidate normalization functions—MinMax, Exponential (Exp), and a combined MinMax+Exp—on the SumMe (Figure 12) and TVSum (Figure 12) datasets.
As expected, MinMax normalization consistently outperformed the other two methods across both segment durations ($W=1,3$) . In particular, shorter segment durations performed better with higher frame consistency weights, while longer segments showed improved performance with lower weights. A similar pattern was observed on the TVSum dataset, where Exp normalization achieved higher performance compared to the others, though the variation between segment durations was less pronounced.
Notably, on both datasets, the combined MinMax+Exp normalization approach yielded competitive results, achieving up to 52% F1 score on SumMe and 60% on TVSum. This highlights the potential of leveraging both normalization strategies to enhance overall performance.
Figure 11: Ablating different normalization strategies and segment durations. F1 scores results on the test set of SumMe dataset across different $\sigma$ values ($\sigma\in[0.0,1.0]$).
Figure 12: Ablating different normalization strategies and segment durations. F1 scores results on the test set of TVSum dataset across different $\sigma$ values ($\sigma\in[0.0,1.0]$).
Emergent Behavior.
Figures 1012 validate the adaptive mechanism encoded in Equation 8—where the scoring function $F_{S}(\text{video})$ dynamically selects the frame-consistency weight $\sigma$ and the segment length $W$ on a per-video basis—and proves essential.
Specifically, we find that a higher frame consistency weight ($\sigma$) leads to better performance when paired with a shorter segment duration ($W$) on an intermediate video duration. This phenomenon is especially pronounced in the TVSum dataset, which is characterized by high-tempo events that span short periods. This highlights the need to assign greater importance to consistency (i.e., higher $\sigma$) while using short segment durations to capture brief, meaningful moments. In contrast, SumMe more often contains shorter, more abstract or low-tempo scenes that unfold over longer durations. In such cases, lower values of $\sigma$ (emphasizing uniqueness over consistency) and longer segments ($W$) are more effective, as discussed previously. Furthermore, for longer videos densely packed with short events, the model tends to select lower $\sigma$ values to emphasize unique, short-lived events, ensuring they are not diluted by surrounding content. These trends confirm that dynamically tuning $\sigma$ and $W$ to each video’s tempo and structure is crucial for producing coherent yet informative summaries across diverse scenarios.
More detailed analysis and ablations. For more ablations and insights we refer the interested reader to Section B where we validate in higher resolution all of the possible configurations.
4.5 Implementation details
We choose to employ frame value density for scene detection, utilizing a dynamic threshold that is defined in section II-A1. This approach allows us to detect scene boundaries effectively by applying the contentDetector function from PySceneDetect  [34], which identifies these boundaries based on the dynamically selected threshold, as shown in II-A1. For feature extraction, we utilize the CLIP/ViT-Large-Patch14 model  [36], which generates frame embeddings for each frame. This model allows us to capture both visual and contextual information from each frame accurately.
For the task of description generation, we utilize the LLAVA/Qwen-2.0-7B Video Language Model [44, 45], which extracts textual information both at the video level and at the scene level. To ensure accurate and contextually rich descriptions, frames are sampled at a rate of 1 frame per second (FPS), with the middle frame of each second selected as the representative input for description generation. This sampling strategy ensures that each frame used captures the essence of its corresponding second, providing a strong basis for textual representation.
As previously noted, processing long videos in their entirety can lead to memory issues. To address this, we employ a batching strategy. Specifically, the sampled frames are divided into manageable batches, with each batch containing 80 frames. Since we sample at 1 FPS, each batch effectively represents 80 seconds of video. This batching approach allows the VideoLM to generate detailed descriptions while maintaining maximum temporal coherence and staying within memory constraints.
We utilize ChatGPT with the GPT-4o [46] model as an LLM judge for scoring scenes based on textual descriptions, setting the temperature parameter to 0.5. The temperature is a key parameter in large language models (LLMs) that controls the randomness or creativity of the generated text. A lower temperature (closer to 0) results in more deterministic and focused outputs, while a higher temperature (closer to 1) leads to more diverse and creative responses. Setting it to 0.5 strikes a balance between coherence and variability, allowing for both accurate and contextually rich descriptions in our summarization process.
Optimizations for Longer Videos. To handle the large number of frames in long videos—particularly in the QFVS dataset, where videos can span up to 5 hours and contain over 220k frames with 45 queries per video—we introduced an optimization strategy. Normally, each video–query pair would be processed individually with the relevant description, which would be computationally expensive and time-consuming, especially when using VideoLM. To reduce overhead, we process all queries at once for each scene. Specifically, instead of evaluating each scene multiple times for different queries, we pass all queries to the LLM in a single prompt and ask it to assign relevance scores to the scene for each query simultaneously. This significantly cuts down on processing time, resources and cost-by up to 50x- though handling long videos remains resource-intensive due to the sheer volume of frames.
All the experiments detailed in this paper were conducted on a PC equipped with one NVIDIA A100-SXM4-40GB GPU and AMD EPYC 7742 64-Core CPUs .
5 Conclusion
  • Training-free state of the art. Our plug-and-play pipeline surpasses all prior unsupervised methods on SumMe and TVSum and rivals the best supervised systems—despite requiring no frame-level annotations, data, domain tuning, or hyper-parameter search.
  • VidLM captions are “summary-ready.” Off-the-shelf VideoLMs already produce rich, coherent scene descriptions that can double as draft summaries. By simply ranking these captions with an LLM judge and propagating scores, we convert them into competitive, user-controllable video skims with no additional training.
  • SOTA LLMs are great text-based “scene judges” With careful prompt design, off-the-shelf LLMs ranked the generated captions in a zero shot manner, leading to SOTA results.
  • Effective query guidance. A single prompt extension lets users steer the summary with natural-language queries; on QFVS we match the fully-supervised CLIP-It while outperforming four classical baselines by up to 20 pp F1.
  • New fine-grained benchmark.
    We release  VidSum-Reason, a new fine-grained text guided video summarization benchmark whose queries demand temporal reasoning and external knowledge, opening a new testbed for more sophisticated video summarization; our method attains a mean 43.4 F1, beating a random baseline by nearly 10 pp.
  • Local trumps global masking. For the video summarization task, measuring the importance of adding a scene to the summary performs better that computing the loss of excluding it from it, this may be explained as descriptions generated from the scene itself outperform those from globally masked scenes withing video videos, confirming the value of focused visual context.
  • CLIP embeddings remain strongest. CLIP/ViT-L14 leads DINO variants by up to 2 pp F1 across datasets in the context of frame embedding for summarization.
  • Match normalization to annotation style. Min–Max works best for SumMe, exponential for TVSum; a hybrid scheme is robust for unseen domains.
  • scalable. Our batch split technique for generating video summaries enables end-to-end processing for a 10-minutes video in realtime and proved to be scalable and robust for a 5-hours video on a single A100 GPU, paving the way for practical deployment.
6 Limitations and feature work
Our work suffers from the following two limitations:
  • Prompt sensitivity. Small wording changes in the LLM judge prompt can modulate scene scores by several points; a formal stability study is still missing.
  • Compute overhead for very long videos. Although processing is real-time for ~10-minute clips, multi-hour footage still requires batched VideoLM inference, and thus the run time is proportional to the number of baches we wish to generate a description for.
Future work induces:
  • Audio–visual fusion. Integrate ASR transcripts and learned audio embeddings so that sound events and speech can influence scene scoring and user queries can reference audio content.
  • Faster summarization. Compress the entire stack to make it faster and enabling it on mobile device, enabling privacy-preserving, real-time highlights on wearable and drones.
Acknowledgment
We acknowledge the use of GPT-4o [46] to assist with grammar correction, typo fixing, and rephrasing of some sentences for improved clarity and readability.
Appendix A Additional implementation details
A.1 LLM Prompt
(a) Determining the optimal threshold using the formulation in Equation 2.
(b) Initial detected scenes based on frame intensity differences, using the optimal threshold defined in Equation 2.
(c) Refined scene boundaries after merging short scenes with their nearest neighbors using cosine similarity over the mean scene embeddings. See Section 3.1.2 for details.
Figure 13: Pipeline flow example (similar to Figure 4) for the Bearpark_climbing video from the SumMe dataset.
"You are tasked with evaluating the importance of a specific scene within a larger video, considering its role in the overall narrative and message of the video. I’ve provided two descriptions below: one for the entire video and one for the specific scene (part) within that video. Your goal is to assess how critical this particular segment is to the understanding or development of the video’s main themes, messages, or emotional impact.
IF user query is available:
User Query: <user query> When assigning a score, consider how well the scene aligns with this preference. Scenes that closely match or contradict the user’s intent should be scored accordingly, reflecting their relevance or irrelevance to the desired summary focus. If the scene is not clearly related to this preference, assign a score based on the default scale and criteria below.
Assign an importance score on a scale of 1 to 100, based on how crucial it is to the overall video. The scale is defined as follows: * 1-20: Minimally important (contributes very little to the overall theme or message) * 21-40: Somewhat important (offers limited context or details that support the main theme) * 41-60: Moderately important (provides useful context or details that support the main theme) * 61-80: Quite important (adds significant context or detail that enhances understanding of the main theme) * 81-100: Highly important (crucial to understanding or conveying the main message of the video)
Video Description: <video description>
Scene Description: <scene description>"
Figure 13 presents an additional pipeline flow example for the Bearpark_climbing video from the SumMe dataset [25]. This example complements the general pipeline in Figure 4, and illustrates how the first stages—Selecting optimal threshold, Initial scene detection, and refined detected scene boundaries—is applied in practice.
A.2 Algorithms implementation
[…]

### v2

@@ B.2 Prompt sensitivity
Prompt Type Dummy Basic Instructive Highly instructive Dummy Basic Instructive Highly instructive
F1-score(%)(raw) 47.8 50.43 49.3 47.7 55.47 56.44 57.78 60.9
F1-score(%)(w/Norm) 43.7 44.5 43.3 47.84 60 60.5 60.62 61.8
F1-score(%)(final) 35.12 43.05 48.84 56.84 59.3 60.8 61.13 62.22
Scene Score Fluctuation(%) Min Mean Max Range Min Mean Max Range
Mean over scenes 16.92 32.46 48.12 31.2 13.22 33.13 54 40.8
Scene score fluctuation under different prompt types. Table 5 reports the impact of prompt design on scene importance scoring for the SumMe and TVSum datasets. We compare four prompt types, ranging from minimal instructions (Dummy) to a fully Highly Instructive prompt that explicitly guides score distribution. The results show that prompt complexity affects both the raw F1-scores and the stability of scene importance assignments. While raw F1-scores may fluctuate across prompt types, performance consistently improves across all stages of our pipeline—from the raw LLM outputs (no normalization or smoothing), through normalization, and finally with normalization and smoothing applied. These gains are most pronounced for the highly instructive prompt, indicating that explicit guidance helps the LLM produce importance scores that better align with the summarization objective. In addition, the table reports scene score fluctuations, which capture the variability of importance scores assigned within each scene. Specifically, Min, Mean, and Max indicate the average minimum, mean, and maximum differences between scene-level scores across the different prompt designs. The Range represents the average total fluctuation for a scene—that is, on average, the score assigned to a scene could vary up to 31.2 points on SumMe dataset and up to 40.8 points on TVSum dataset across its scenes.
(a) Average processing time for SumMe, TVSum and VidSum-Reason videos as a function of frame count. A linear trend is observed, with slope differences explained by varying event distributions across datasets.
(b) Processing time for QFVS videos at three different sampling rates. Lower FPS yields shorter processing times while preserving dataset coverage, with a near-linear scaling indicated by the fitted line.
Figure 16: Comparison of processing time trends across datasets and sampling strategies. Figure 16(a) presents the average processing time for short videos (SumMe, TVSum and VidSum-Reason) against frame count, showing a clear linear relationship. Figure 16(b) illustrates the processing time for QFVS long videos under different frame sampling rates. Lower FPS significantly reduces computational cost while retaining full content coverage.
Appendix C Our pipeline’s processing time
To evaluate the computational efficiency of our approach, we measured the average processing time for each dataset, alongside the corresponding dataset size and average frame count (Table 6). For smaller-scale datasets such as SumMe and TVSum, the complete pipeline (including frame extraction, scene detection, description generation, scene and frame scoring) executes within approximately 12–15 minutes per video on average. Larger datasets, particularly  VidSum‑Reason, require slightly longer processing times due to their higher frame counts and the exiting of user provided queries.
Figure 16(a) shows the average processing time for videos from SumMe, TVSum and VidSum-Reason as a function of frame count. A clear linear trend is observed, indicating that processing time scales proportionally with the number of frames. Differences in slope across datasets can be attributed to variations in event distribution: SumMe videos are less densely packed with events compared to TVSum and VidSum‑Reason, resulting in differences in scene complexity, scene detection, and description generation times. Despite these variations, the overall relationship remains consistent, highlighting the importance of strategies such as frame-rate reduction to maintain scalability for longer videos.
To address scalability challenges when processing longer videos—particularly the high computational cost of the description generation stage—we adopted a lower frame rate sampling strategy for long-form videos. Specifically, for the QFVS dataset, we sampled videos at reduced FPS rates (0.5 and 0.2) compared to the default 1.0 FPS.
The fitted trend line for short videos confirms a near-linear relationship between frame count and processing time (see Figure 16(b)). In Table 6, we report the average processing time for each dataset and the percentage of time spent on each stage of our pipeline.
As expected, long videos (QFVS videos) required significantly more time, with the description generation stage accounting for on average 52% of the total processing time. To reduce this bottleneck, we evaluated our pipeline at lower FPS rates (0.5 and 0.2). While this reduced the description generation time substantially (from 317 to 164 and 112 minutes for 1.0, 0.5, and 0.2 FPS, respectively), the F1-score decreased by approximately 2 percentage points. This decrease is an expected effect of lower frame sampling, which directly impacts the quality of the generated descriptions and subsequent LLM judgments.
We note that the reported runtime in Table 6 corresponds to a sequential batch-wise description generation protocol. This step can be naturally embarrassingly parallel across batches (e.g., via multiple concurrent calls), and thus the wall-clock time can be reduced proportionally with the degree of parallelization.
Table 6: Processing time and frame count across datasets. For datasets with multiple queries per video ( VidSum‑Reason and QFVS), the average number of queries per video is reported in parentheses.
Dataset #Videos (#Query) $\mathbf{\overline{\#Frames}}$ $\mathbf{\overline{Time\,(min)}}$ $\mathbf{\overline{Scene\,detection\,(\%)}}$ $\mathbf{\overline{Description\,generation\,(\%)}}$ $\mathbf{\overline{LLM\,scoring\,(\%)}}$ $\mathbf{\overline{Frame\,weighting\,(\%)}}$ F1-score(%)
SumMe 25 (0) 4,393 12 0.54 0.35 0.01 0.13 56.84
TVSum 50 (0) 7,047 15.2 0.44 0.38 0.02 0.18 62.22
vidSum-Reason 9 (2.22) 7,872 20.7 0.55 0.20 0.00 0.19 43.40
QFVS (1.0 fps) 4 (45) 228,994 610 0.23 0.52 0.08 0.08 53.42
QFVS (0.5 fps) 4 (45) 228,994 530 0.44 0.31 0.08 0.09 52.60
QFVS (0.2 fps) 4 (45) 228,994 470 0.47 0.24 0.09 0.10 51.38
Appendix D Additional details about our dataset: VidSum‑Reason
In Table 8, we provide the full  VidSum‑Reason dataset. Each video-query pair is accompanied by its query class and the corresponding textual query. Additionally, Figure 17 shows the frequency of each query class in the VidSum‑Reason dataset.
Additionally, we provide metadata for our dataset (see Table  D), VidSum‑Reason, including FPS, duration, number of frames, and resolution. The original videos can be accessed via their URL as follows: https://www.youtube.com/watch?v=<video ID>.
Table 7: VidSum‑Reason dataset metadata. Video-level statistics of the dataset, reporting frame rate, temporal duration, total frame count, and spatial resolution.
Video ID FPS Duration (s) #Frames Resolution
[…]
Average 30.70 240.24 7804
Figure 17: VidSum‑Reason: Frequency distribution of query categories.
Table 8: List of query-annotated videos in the  VidSum‑Reason dataset. Each row corresponds to a video, showing the video ID, the query class (which defines the type of information requested), and the user-provided query used to guide summarization. This table highlights the diversity of query intents and supports evaluation of personalized video summarization.
Video ID Query Category Text Query GIVerZ9mUpU Reasoning with General Knowledge Filter any scenes that involve violence wvDENCN4i3c Reasoning Focus on emotional reactions and scenes 5L4DQfVIcdg Standard with Special Attribute Focus on scenes with the coin machine 5L4DQfVIcdg Reasoning Highlight scenes showing the passage of time 5L4DQfVIcdg Reasoning Prioritize scenes that demonstrate emotions 3XA0bB79oGc Standard Focus on scenes with both the boy and the puppy 3XA0bB79oGc Reasoning Exclude scenes with negative emotions 3XA0bB79oGc Reasoning Prioritize scenes with emotional shifts 77AA5bWCGlE Reasoning with General Knowledge Focus on scenes where the score changes 77AA5bWCGlE Standard with Special Attribute Highlight the free throws 5dY1kU_3fTI Reasoning Highlight every scene where measurement is involved 5dY1kU_3fTI Standard with Special Attribute Highlight every scene where non-power (hand) tools are used GgtUWB405Mo Standard with Special Attribute Focus on cars with bright colors GgtUWB405Mo Reasoning with General Knowledge Highlight crossover vehicles in the video GgtUWB405Mo Reasoning with General Knowledge Prioritize German cars EXeTwQWrcwY Standard Show scenes with explosions EXeTwQWrcwY Reasoning with General Knowledge Show scenes that feature the Joker or that feel related to him EXeTwQWrcwY Reasoning Highlight moments that seem like public spectacle or protest 9QABNfDhSxs Reasoning with General Knowledge Exclude scenes that aren’t PG-13 appropriate 9QABNfDhSxs Standard Focus on scenes that show blood
Annotation process.
VidSum‑Reason was created as a proof of concept to demonstrate that LLM–VideoLM systems, when properly orchestrated, can perform robust video summarization for long-tailed and reasoning-heavy queries beyond surface-level prompts.
The dataset was annotated by a single primary annotator, followed by a structured verification process conducted by two additional reviewers. Because these reviewers performed quality control rather than independent re-annotation, standard inter-annotator agreement metrics (e.g., Cohen’s Kappa), which require multiple independent labels per item, are not applicable under their conventional definition.
[…]